# footnotetext: MSC2020: 53A04, 53A70 (primary), 53A17, 70B15, 37K25, 37K10, 35Q53 (secondary)

### v2

@@ 3. Curves and integrable systems
There are deep connections between differential geometry and the theory of integrable systems, where the integrable differential or difference equations arise as compatibility conditions of the geometric objects (see, for example, [27]). In the context of the deformation of spatial and plane curves, the Frenet frame and its deformation are described by the system of linear partial differential equations. The compatibility condition yields so-called the Ablowitz-Kaup-Newell-Segur (AKNS) hierarchy for the curvature and the torsion, including the nonlinear Schrödinger (NLS) equation as a typical example. In particular, if the torsion is a constant, the integrable deformation is governed by the mKdV equation and its hierarchy [9, 23, 27]. The connections are also investigated in a discrete setting, where curves are discretised as piecewise linear (polygonal) curves. For example, continuous deformations of discrete plane and spatial curves have been considered in [5], where the deformations are governed by the semi-discrete analogue of the mKdV and the NLS equations. Discrete deformations of the discrete plane and spatial curves have been studied in [16]. In principle, for a given solution to the nonlinear differential/difference equation for the curvature and torsion, one can reconstruct the curve by solving the associated system of linear differential/difference equations for the Frenet frame and integrating the tangent vector.
The theta functions naturally appear in so-called the quasi-periodic solutions to the nonlinear integrable equations, and the corresponding solutions to the associated linear problem are nothing but the Baker-Akhiezer functions. Reconstruction of the deformation of the spatial curve with constant torsion described by the theta function solutions to the mKdV equation has been studied in this framework in [4, 24]. Also, an explicit formula of the spatial curves in terms of the $\tau$ functions of the two-component KP hierarchy based on Hirota’s bilinear formalism has been presented in [11].
Since the dynamics of the Kaleidocycles are modelled by the integrable deformation of the closed discrete spatial curves with a constant torsion angle which preserves the segment-length and torsion angle [19], we study the theta function solutions to the corresponding integrable semi-discrete equations, where the theta functions are nothing but the $\tau$ functions. Therefore we construct the explicit formula of the Kaleidocycles based on the method of direct reconstruction of the discrete spatial curves in terms of the $\tau$ functions.
Hereafter, a (discrete) curve always indicates a polygonal spatial curve.
@@ 3.1. Expression of a curve by the $\tau$ functions
We construct a discrete curve $\gamma_{n}$ and its Frenet frame $\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ indexed by $n\in\mathbb{Z}$ using a modification of the $\tau$ functions presented in [11]. Let $F_{n},\widetilde{G}_{n}:\mathbb{C}\to\mathbb{R}$ and $f_{n},g_{n},G_{n},H_{n}:\mathbb{C}\to\mathbb{C}$ be functions indexed by $n\in\mathbb{Z}$ with a complex variable $z$ that satisfy the following equations
(3.1) $f_{n}f_{n}^{*}+g_{n}g_{n}^{*}=F_{n+1}F_{n}$
(3.2) $\beta F_{n}f_{n}+G_{n}g_{n}^{*}=F_{n+1}f_{n-1}$
(3.3) $\beta F_{n}g_{n}-G_{n}f_{n}^{*}=F_{n+1}g_{n-1}$
(3.4) $H_{n+1}F_{n}-H_{n}F_{n+1}=f_{n}^{*}g_{n}$
(3.5) $D_{z}F_{n+1}\cdot F_{n}:=\frac{dF_{n+1}}{dz}F_{n}-F_{n+1}\frac{dF_{n}}{dz}=g_{n}g_{n}^{*}$
@@
where $\beta\in\mathbb{R}$ is some constant and $D_{z}$ is Hirota’s bilinear differential operator [12]. These functions are called the $\tau$ functions. Let $\varrho_{n}=\dfrac{1}{2}\dfrac{G_{n}}{\widetilde{G}_{n}}$ and
(3.10) $T_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vspace{2mm}f_{n}^{*}g_{n}+f_{n}g_{n}^{*}\\ \vspace{2mm}\frac{1}{i}\left(f_{n}^{*}g_{n}-f_{n}g_{n}^{*}\right)\\ f_{n}f_{n}^{*}-g_{n}g_{n}^{*}\end{array}\right)\right|_{z=0}$
(3.14) $\widetilde{N}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vspace{2mm}\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \vspace{2mm}\frac{1}{i}\left[\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}-\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ -2\left(f_{n}^{*}g_{n}^{*}\varrho_{n}+f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0}$
(3.18) $\widetilde{B}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vspace{2mm}\frac{1}{i}\left[-\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ \vspace{2mm}\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \frac{2}{i}\left(f_{n}^{*}g_{n}^{*}\varrho_{n}-f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0}$
(3.22) $\gamma_{n}$ $=\left.\left(\begin{array}[]{c}\vspace{2mm}\dfrac{H_{n}+H_{n}^{*}}{F_{n}}\\ \vspace{2mm}\dfrac{1}{i}\dfrac{H_{n}-H_{n}^{*}}{F_{n}}\\ n-2\dfrac{\partial\log F_{n}}{\partial z}\end{array}\right)\right|_{z=0}.$
Similarly to [11, §3], we obtain:
@@
$\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ defines the discrete Frenet frame for the curve $\gamma$ satisfying (2.6). Its curvature and torsion angles are given by
(3.23) $\lambda_{n}$ $=\left.\frac{1}{i}\log\left(\frac{G_{n}^{*}G_{n-1}}{\widetilde{G}_{n}\widetilde{G}_{n-1}}\right)\right|_{z=0}=\left.\frac{1}{2i}\log\left(\frac{G_{n}^{*}G_{n-1}}{G_{n}G_{n-1}^{*}}\right)\right|_{z=0},$
$\kappa_{n}$ $=\left.2\arctan\left(\frac{\widetilde{G}_{n}}{\beta F_{n}}\right)\right|_{z=0}.$
3.2. Motion of Kaleidocycle on the osculating plane
@@
where $\tilde{\rho}>0$ is a constant and $\dot{\gamma}_{n}=\dfrac{d\gamma_{n}}{dt}$. These conditions mean that the velocity of each vertex is restricted within the osculating plane and has a constant magnitude. It is shown in [19] that the Frenet-Serret formula (2.6) together with (3.24) demand that there exists a function $\Theta_{n}$, called the potential function, with $\kappa_{n}=\dfrac{\Theta_{n+1}-\Theta_{n-1}}{2}$ that satisfies either the semi-discrete potential mKdV equation:
(3.25) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\left(1+\cos\lambda\right)\tilde{\rho}\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
or the semi-discrete sine-Gordon equation
(3.26) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\left(1-\cos\lambda\right)\tilde{\rho}\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right).$
The motions corresponding to these equations differ in general. However, the explicit construction discussed in §5 satisfies both equations simultaneously.
@@
As noted in §2.1, the dimension of the configurations of a Kaleidocycle defined by (2.2) is generally $k-6$, and hence, (3.24) may look too restrictive. However, the Möbius Kaleidocycles, which exhibit critical torsion angles, the dimension of the configurations is conjectured to be one regardless of $k$. An extensive numerical simulation suggests that the motion of the Möbius Kaleidocycle satisfies (2.2). Therefore, (2.2) specifies a particular yet essential motion of Kaleidocycles.
4. An explicit formula of closed discrete curve with a constant torsion angle in terms of elliptic theta functions
Based on the representation of curves discussed in §3.1, we define certain $\tau$ functions satisfying (3.1)-(3.6) using elliptic theta functions to construct a family of discrete curves with a constant torsion angle. Then, we give sufficient conditions for the curve to be closed and derive an explicit expression for the motion of Kaleidocycles.
4.1. A family of open curves with a constant torsion angle
@@
(4.1) $\vartheta_{1}(v;w)$ $=-\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)\left(v+\frac{1}{2}\right)},$
$\vartheta_{2}(v;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)v},$
$\vartheta_{3}(v;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi inv},$
$\vartheta_{4}(v;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi in\left(v+\frac{1}{2}\right)}.$
We often denote $\vartheta_{j}(v;w)$ by $\vartheta_{j}(v)$ when $w$ is fixed.
We construct the $\tau$ functions satisfying (3.1)-(3.6) using these elliptic theta functions such that the torsion angle given by Proposition 3.1 is a constant. The following observations are useful insights for constructing solutions. In order to make $F$ a real-valued function, it is convenient to use a property of elliptic theta functions, shown in Lemma K.1. Therefore, $w$ is taken to be a pure imaginary number. Additionally, to make $\gamma$ a closed curve, it is convenient to use the pseudo-period of elliptic theta functions (A.12), and therefore the space variable is taken to be a pure imaginary number. A real parameter $r$ is introduced to prevent the curve from being a planar curve. Then, the solution to satisfy (3.1)-(3.6) can be constructed from the identities of elliptic theta functions (e.g. (A.2)), and the result is as shown in Theorem 4.1. For $x\in\mathbb{C}$ and $y>0$, we define $d_{j}(x;y)=\frac{\vartheta^{{}^{\prime}}_{j}(x;iy)}{\vartheta_{j}(x;iy)}$, where ${}^{{}^{\prime}}$ denotes the derivative with respect to the first variable of elliptic theta functions. For $y>0$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, we define
(4.2) $\alpha_{2}=\sqrt{\vartheta_{3}(2r;iy)\vartheta_{3}(0;iy)},\quad\beta=\frac{\vartheta_{3}(iv;iy)}{\vartheta_{3}(0;iy)},\quad s_{2}=\vartheta_{3}(2r;iy)\vartheta_{1}(iv;iy),$
$u_{2}=\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right),$
$\Delta_{j}=d_{j}\left(-\frac{1}{2}iv+r;y\right)-d_{j}\left(\frac{1}{2}iv+r;y\right),$
$R_{j}=\frac{\vartheta_{j}\left(-\frac{1}{2}iv+r;iy\right)}{\vartheta_{j}\left(\frac{1}{2}iv+r;iy\right)}.$
Noting $\vartheta_{j}(z;iy)^{*}=\vartheta_{j}(z^{*};iy)$ when $y>0$, we have
(4.3) $\begin{split}&\alpha_{2}^{*}=\alpha_{2},\quad\beta^{*}=\beta,\quad s_{2}^{*}=-s_{2},\quad u_{2}^{*}=u_{2},\\ &\Delta_{j}^{*}=-\Delta_{j},\quad R_{j}R_{j}^{*}=1.\end{split}$
@@
(4.4) $F_{n}(z)=\alpha_{2}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv\right),$
$f_{n}(z)=\vartheta_{3}\left(-\frac{1}{2}iv+r\right)R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}+r\right),$
$g_{n}(z)=\vartheta_{1}\left(\frac{1}{2}iv+r\right)R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-r\right),$
$G_{n}(z)=\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right),$
$H_{n}(z)=\frac{u_{2}}{\alpha_{2}s_{2}}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r\right),$
$\widetilde{G}_{n}(z)=-i\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right).$
Then, they satisfy (3.1)-(3.6) to form $\tau$ functions. The corresponding discrete curve $\gamma$ defined by (3.10) has a unit segment length $\left|\gamma_{n+1}-\gamma_{n}\right|=1$. The torsion angle is given by
(4.5) $\lambda_{n}=\frac{1}{i}\log\left(\frac{G_{n}^{*}(0)G_{n-1}(0)}{\widetilde{G}_{n}(0)\widetilde{G}_{n-1}(0)}\right)=\frac{1}{i}\log\left(R_{1}R_{3}^{-1}\right).$

### v3

@@ 3. Curves and integrable systems
There are deep connections between differential geometry and the theory of integrable systems, where the integrable differential or difference equations arise as compatibility conditions of the geometric objects (see, for example, [29]). In the context of the deformation of spatial and plane curves, the Frenet frame and its deformation are described by the system of linear partial differential equations. The compatibility condition yields the so-called Ablowitz-Kaup-Newell-Segur (AKNS) hierarchy for the curvature and the torsion, including the nonlinear Schrödinger (NLS) equation as a typical example. In particular, if the torsion is a constant, the integrable deformation is governed by the mKdV equation and its hierarchy [10, 25, 29]. The connections are also investigated in a discrete setting, where curves are discretised as piecewise linear (polygonal) curves. For example, continuous deformations of discrete plane and spatial curves have been considered in [6], where the deformations are governed by the semi-discrete analogue of the mKdV and the NLS equations. Discrete deformations of the discrete plane and spatial curves have been studied in [18]. In principle, for a given solution to the nonlinear differential/difference equation for the curvature and torsion, one can reconstruct the curve by solving the associated system of linear differential/difference equations for the Frenet frame and integrating the tangent vector.
The theta functions naturally appear in so-called quasi-periodic solutions to the nonlinear integrable equations, and the corresponding solutions to the associated linear problem are nothing but the Baker-Akhiezer functions. Reconstruction of the deformation of the spatial curve with constant torsion described by the theta function solutions to the mKdV equation has been studied in this framework in [4, 26]. Also, an explicit formula of the spatial curves in terms of the $\tau$ functions of the two-component KP hierarchy based on Hirota’s bilinear formalism has been presented in [12].
@@ 3.1. Expression of a curve by the $\tau$ functions
We introduce a special class of discrete curves $\gamma_{n}$ and their Frenet frames $\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ indexed by $n\in\mathbb{Z}$ using a modification of the $\tau$ functions presented in [12]. Let $F_{n},\widetilde{G}_{n}:\mathbb{C}\to\mathbb{R}$ and $f_{n},g_{n},G_{n},H_{n}:\mathbb{C}\to\mathbb{C}$ be functions indexed by $n\in\mathbb{Z}$ with a complex variable $z$ that satisfy the following equations
(3.1) $f_{n}f_{n}^{*}+g_{n}g_{n}^{*}=F_{n+1}F_{n},$
(3.2) $\beta_{1}F_{n}f_{n}+G_{n}g_{n}^{*}=F_{n+1}f_{n-1},$
(3.3) $\beta_{1}F_{n}g_{n}-G_{n}f_{n}^{*}=F_{n+1}g_{n-1},$
(3.4) $H_{n+1}F_{n}-H_{n}F_{n+1}=f_{n}^{*}g_{n},$
(3.5) $D_{z}F_{n+1}\cdot F_{n}:=\frac{dF_{n+1}}{dz}F_{n}-F_{n+1}\frac{dF_{n}}{dz}=g_{n}g_{n}^{*},$
@@
where $\beta_{1}\in\mathbb{R}$ is some constant and $D_{z}$ is Hirota’s bilinear differential operator [13], and the symbol * denotes the complex conjugate. These functions are called the $\tau$ functions. Let $\varrho_{n}=\dfrac{1}{2}\dfrac{G_{n}}{\widetilde{G}_{n}}$ and
(3.10) $T_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vskip 5.69054ptf_{n}^{*}g_{n}+f_{n}g_{n}^{*}\\ \vskip 5.69054pt\frac{1}{i}\left(f_{n}^{*}g_{n}-f_{n}g_{n}^{*}\right)\\ f_{n}f_{n}^{*}-g_{n}g_{n}^{*}\end{array}\right)\right|_{z=0},$
(3.14) $\widetilde{N}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vskip 5.69054pt\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \vskip 5.69054pt\frac{1}{i}\left[\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}-\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ -2\left(f_{n}^{*}g_{n}^{*}\varrho_{n}+f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0},$
(3.18) $\widetilde{B}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vskip 5.69054pt\frac{1}{i}\left[-\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ \vskip 5.69054pt\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \frac{2}{i}\left(f_{n}^{*}g_{n}^{*}\varrho_{n}-f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0},$
(3.22) $\gamma_{n}$ $=\left.\left(\begin{array}[]{c}\vskip 5.69054pt\dfrac{H_{n}+H_{n}^{*}}{F_{n}}\\ \vskip 5.69054pt\dfrac{1}{i}\dfrac{H_{n}-H_{n}^{*}}{F_{n}}\\ n-2\dfrac{\partial\log F_{n}}{\partial z}\end{array}\right)\right|_{z=0}.$
Similarly to [12, §3], we obtain:
@@
$\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ defines the discrete Frenet frame for the curve $\gamma$ satisfying the discrete Frenet-Serret formula (2.5). Its curvature and torsion angles are given by
(3.23) $\lambda_{n}$ $=\left.\arccos\left(\frac{G_{n}^{*}G_{n-1}+G_{n}G_{n-1}^{*}}{2\widetilde{G}_{n}\widetilde{G}_{n-1}}\right)\right|_{z=0},$
$\kappa_{n}$ $=\left.2\arctan\left(\frac{\widetilde{G}_{n}}{\beta_{1}F_{n}}\right)\right|_{z=0},$
where we take the principal value for $\arccos$ in $0,\pi]$ and for $\arctan$ in $(-\pi/2,\pi/2)$.
3.2. Motion of a Kaleidocycle on the osculating plane
@@
where $\rho\neq 0$ is a constant and $\dot{\gamma}_{n}=\dfrac{d\gamma_{n}}{dt}$. These conditions imply that the velocity of each vertex is restricted to the osculating plane and has a constant magnitude. It is shown in [21] that the Frenet-Serret formula (2.5), together with (3.24), requires that there exists a function $\Theta_{n}$, called the potential function, with $\kappa_{n}=\dfrac{\Theta_{n+1}-\Theta_{n-1}}{2}$ that satisfies either the semi-discrete potential mKdV equation:
(3.25) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\left(1+\cos\lambda\right)\rho\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
or the semi-discrete sine-Gordon equation:
(3.26) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\left(1-\cos\lambda\right)\rho\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right).$
The motions corresponding to these equations differ in general. However, we will construct a motion in §5 that satisfies both equations simultaneously.
@@
As noted in §2.1, the dimension of the configurations of a Kaleidocycle defined by (1.1) is generally k6k-6. However, for the Möbius Kaleidocycles, which exhibit critical torsion angles, the dimension of the configurations is conjectured to be one, regardless of $k$. An extensive numerical simulation suggests that the motion of the Möbius Kaleidocycle satisfies (3.24). Therefore, (3.24) specifies a particular yet essential motion of Kaleidocycles.
4. An explicit formula of a discrete curve with a constant torsion angle in terms of elliptic theta functions
Based on the representation of curves discussed in §3.1, we define certain $\tau$ functions satisfying (3.1)-(3.6) using elliptic theta functions to construct a family of discrete curves with a constant torsion angle in Theorem 4.2.
4.1. A family of curves with a constant torsion angle
@@
(4.1) $\vartheta_{1}(u;w)$ $=-\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)\left(u+\frac{1}{2}\right)},$
$\vartheta_{2}(u;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)u},$
$\vartheta_{3}(u;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi inu},$
$\vartheta_{4}(u;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi in\left(u+\frac{1}{2}\right)}.$
We often denote $\vartheta_{j}(u;w)$ by $\vartheta_{j}(u)$ when $w$ is fixed. These functions satisfy various identities, collected in Appendix A, that are convenient to construct the $\tau$ functions satisfying (3.1)–(3.6). We see the torsion angle given by Proposition 3.1 is a constant.
The following property is useful for examining whether a $\tau$ function constructed in this paper is real-valued.
Lemma 4.1.
The following holds for any $u\in\mathbb{C}$ when $w=iy$ $(y>0)$.
(4.2) $\vartheta_{j}(u)^{*}=\vartheta_{j}(u^{*}).$
Proof.
Consider the following theta function with characteristics.
(4.3) $\vartheta_{a,b}(u):=\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(n+a)(u+b)+\pi i(n+a)^{2}w\right),\quad a,b\in\left\{0,\frac{1}{2}\right\},$
then,
(4.4) $\vartheta_{1}(u)=-\vartheta_{\frac{1}{2},\frac{1}{2}}(u),\quad\vartheta_{2}(u)=\vartheta_{\frac{1}{2},0}(u),\quad\vartheta_{3}(u)=\vartheta_{0,0}(u),\quad\vartheta_{4}(u)=\vartheta_{0,\frac{1}{2}}(u),$
hold. Since $w^{*}=-w$, we have
(4.5) $\begin{split}\vartheta_{a,b}(u)^{*}=&\sum_{n\in\mathbb{Z}}\exp\left(-2\pi i(n+a)(u^{*}+b)-\pi i(n+a)^{2}w^{*}\right)\\ =&\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(-n-a)(u^{*}+b)+\pi i(n+a)^{2}w\right)\\ =&\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(n+a)(u^{*}+b)+\pi i(n+a)^{2}w\right)\\ =&\vartheta_{a,b}(u^{*}).\end{split}$
For $u\in\mathbb{C}$ and $y>0$, we define $d_{j}(u;y)=\frac{\vartheta^{{}^{\prime}}_{j}(u;iy)}{\vartheta_{j}(u;iy)}$, where $^{{}^{\prime}}$ denotes the derivative with respect to the first variable of elliptic theta functions. For $y>0$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, we define
(4.6) $u_{1}=\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right),$
$\Delta_{j}=d_{j}\left(-\frac{1}{2}iv+r;y\right)-d_{j}\left(\frac{1}{2}iv+r;y\right),\quad R_{j}=\frac{\vartheta_{j}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{j}\left(\frac{1}{2}iv+r\right)},$
where $\vartheta_{j}(u;iy)$ is denoted by $\vartheta_{j}(u)$. Using Lemma4.1, we see
(4.7) $\begin{split}u_{1}>0,\quad\Delta_{j}^{*}=-\Delta_{j},\quad R_{j}^{*}=R_{j}^{-1}.\end{split}$
Note that from (4.7), we see $R_{1}R_{3}^{*}=R_{1}R_{3}^{-1}$ and is a complex number of magnitude $1$. Thus
(4.8) $\begin{split}\frac{R_{1}R_{3}^{*}+(R_{1}R_{3}^{*})^{*}}{2}=\frac{R_{1}R_{3}^{-1}+R_{1}^{-1}R_{3}}{2},\end{split}$
is real and lies in [1,1][-1,1]. Also, from (4.7) and
(4.9) $\Delta_{3}-\Delta_{1}=-\frac{\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{1}^{{}^{\prime}}\left(0\right)}{u_{1}},$
we see that $\Delta_{3}-\Delta_{1}\neq 0$.
@@
(4.10) $F_{n}(z)=\alpha_{1}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv\right),$
$f_{n}(z)=\alpha_{2}R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}+r\right),$
$g_{n}(z)=\alpha_{3}R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-r\right),$
$G_{n}(z)=\alpha_{4}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right),$
$H_{n}(z)=\alpha_{5}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r\right),$
$\widetilde{G}_{n}(z)=-i\alpha_{4}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right),$
where
$\alpha_{1}=\sqrt{\vartheta_{3}(2r)\vartheta_{3}(0)},\quad\alpha_{2}=\vartheta_{3}\left(-\frac{1}{2}iv+r\right),\quad\alpha_{3}=\vartheta_{1}\left(\frac{1}{2}iv+r\right),$
$\alpha_{4}=\frac{\vartheta_{1}(iv)\sqrt{\vartheta_{3}(2r)\vartheta_{3}(0)}}{\vartheta_{3}(0)},\quad\alpha_{5}=\frac{u_{1}}{\vartheta_{3}(2r)\vartheta_{1}(iv)\sqrt{\vartheta_{3}(2r)\vartheta_{3}(0)}}.$
Then, they satisfy (3.1)-(3.6) with
(4.11) $\beta_{1}=\frac{\vartheta_{3}(iv)}{\vartheta_{3}(0)},$
and form $\tau$ functions. The corresponding discrete curve $\gamma$ defined by (3.10) has a unit segment length $\left|\gamma_{n+1}-\gamma_{n}\right|=1$. The torsion angle is given by
(4.12) $\begin{split}\lambda_{n}&=\arccos\left(\frac{G_{n}^{*}(0)G_{n-1}(0)+G_{n}(0)G_{n-1}(0)^{*}}{2\widetilde{G}_{n}(0)\widetilde{G}_{n-1}(0)}\right)\\ &=\arccos\left(\frac{R_{1}R_{3}^{-1}+R_{1}^{-1}R_{3}}{2}\right).\end{split}$